OpenStudy (anonymous):

Solve for x 67=0.4e^5x

OpenStudy (anonymous):

okay, this one is easy

OpenStudy (anonymous):

ok where do I begin

OpenStudy (anonymous):

Firstly, divide both sides by 0.4 so that we can cancel it out

OpenStudy (anonymous):

ok \[\frac{ 67 }{ 0.4}\]=e^5x

OpenStudy (anonymous):

Do you know what you get if you take the natural logarithm of e?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Try it in your calculator

OpenStudy (anonymous):

I got 1

OpenStudy (anonymous):

\[\ln(\frac{67}{0.4})=5x*\ln(e)\]

OpenStudy (anonymous):

perfect, so we know that ln(e)=1, do you see how to solve the rest of the problem?

OpenStudy (anonymous):

I think so hold on a minute and let me do it

OpenStudy (anonymous):

so is the answer 1.024

OpenStudy (anonymous):

It's easy to check, just plug it right back into the equation

OpenStudy (anonymous):

yup ok thanks

OpenStudy (anonymous):

Did you have any more logarithm questions? also, any question involving something like 0.4e, if you can divide out the coefficient, you will always know that ln(e)=1, which makes the rest a lot easier for you.

OpenStudy (anonymous):

I do it is Solve for x again [4e ^{x}+e ^{x}(4x-3)=0\]

OpenStudy (anonymous):

I'm not sure how that's formatted